OpenStudy (anonymous):

What is the vertex of the graph of g(x) = |x - 9| - 3?

OpenStudy (anonymous):

is it (9,-3)... y=g(x) = |x - 9| - 3 or y= |x - 9| - 3 or y+3== |x - 9| let x-9=X and y+3=Y thus the graph reduces to Y=|X| whose vertex is (0,0) (wrt (X,Y)) thus putting x-9=0 and y+3=0 gives the reqd vertex as (9,-3)